Exercise 16-5

Following is the balance sheet of the BDO Partnership:

Cash $12,000 Liabilities $12,000
Accounts Receivable 41,000 Brink, Capital 47,000
Inventory 28,000 Davis, Capital 25,000
Equipment 58,000 Olsen, Capital 55,000
$139,000 $139,000


The partners share income 40:40:20, respectively. Assume that 70% of the receivables are collected and that inventory with a book value of $14,000 is sold for $10,000. All cash available at this time is to be distributed.

Determine the proper distribution of cash, using the safe payment approach. (Round answers to 0 decimal places, e.g. 5,125. Enter credit balance of an account and credit posting to an account with negative sign preceding the number, e.g. -45 or parentheses, e.g. (45).)

Capital Balances
Cash Noncash
Assets
Liabilities Brink
40%
Davis
40%
Olsen
20%

Account BalancesPayment to CreditorsPayment to PartnersSale of Inventory, Collect Accounts Receivable

$

$

$

$

$

$

Account BalancesPayment to CreditorsPayment to PartnersSale of Inventory, Collect Accounts Receivable

Account BalancesPayment to CreditorsPayment to PartnersSale of Inventory, Collect Accounts Receivable

Account BalancesPayment to CreditorsPayment to PartnersSale of Inventory, Collect Accounts Receivable

$

$

$

$

$

$
Answer #1

Given ,

The profit sharing ratio of partners is 40:40:20.

Cash balance after adjustments:

Opening balance = 12,000

Cash received from sale of inventory = 10,000

Cash collected from accounts receivables = 41,000*70% = 28,700

Cash balance after above adjustments = 50,700

Non cash assets:

Balance of account receivable = 41,000*30% = 12,300

Balance of inventory = 28,000-14000 = 14000

Equipment = 58,000

Total non cash assets = 84,300

Liabilities: = 12,000

Capital Balances :

Brink Davis Olsen
Opening balance 47000 25000 55000
Cash distributed 20280 20280 10140
Closing balance 67280 45280 65140